Determine the power dissipated when a current of 10Ma flows through an appliance having a resistance of 8kΩ
1
Expert's answer
2020-04-30T10:42:42-0400

As per the given question,

Current in the appliances "(i)=10\times 10^6A"

Resistance of the appliances "(R)=8\times 10^3 \Omega"

We know that, power dissipation in the appliances "(P)=i^2R"

"P=(10\times 10^6)^2\times8000=8\times 10^{17}watt"